Find the solution to the given systems

(a) X'=\begin{bmatrix} 1/2 &0 \\ 1& -1/2 \end{bmatrix}X X(0)=\bigl(\begin{smallmatrix} 3\\ 5 \end{smallmatrix}\bigr)

(b) dx/dt=5x+y dy/dt=-2x+3y
Answer #1

given X Xlo) +112 Here A = 12 12 Lel Abe an eigenvalue of A then det (A-AI)=0 72-7 = (3-1)(31-4) = 0 -1 2-2 7 = ² 1 - 1 / 212 = -1 then (A + 2 IX [***]]-1 no Mence X2 = 312 82 72 let na = 1 is an eigenvector of A is then X2 = (0) A corresponding to- Given dx Sxty f dy at = -2X+34 - dt Differentiating o o.sito l get x = sx ty > X = Sx + (-2X+34) using y= -2X+34 from ②44 tett 4 et Cc, cost & Casint) (-c, sind + C₂ cost) = sett (a cosa tez sint ) + x Y = -cte (ci cose + C2 sind) + All-C, sint